# M2Chat: Empowering VLM for Multimodal LLM Interleaved Text-Image Generation

### v2

@@ 3.1 Preliminary
We begin by detailing our instruction design process. We draw from an image-text dataset $\mathcal{D}:\{\mathcal{X,Y}\}$, containing pairs of images $\{x\}_{i=1}^{N}$ and their corresponding textual contexts $\{y\}_{i=1}^{N}$, where $N$ is the sample count. To construct the context $Y$, we adopt the principles of visual instruction tuning [21] and introduce an additional image token $<|img|>$ to denote padding, alongside $<|IC|>$ to signal the start of an image caption. These tokens serve as markers to differentiate token types during the two-stage M3FT training phase. The context $y$ is organized as
@@ 3.2 Framework Architecture
where $l$ indicates the $l^{th}$ layer. However, aligning features from two modalities directly limits model creativity, particularly in multimodal conversation. Therefore, we introduce a multi-level feature fusion strategy that incorporates granular low-level visual primitive $f_{img}$ encoding layouts and textures into semantically rich deep high-level multimodal feature $h_{l}$ to obtain the final fused feature $f_{fus}$. A learnable gate modulates this fusion for adaptability. This method enriches image generation guidance, tailoring output to varied task demands. We differentiate answer tokens $t_{ans}$ and caption tokens $t_{cap}$ from VLM output $h_{l}$ with special tokens $<|IC|>$, computing their cosine similarity, which acts as a task-specific flag signaling when answers and captions diverge or require consistency in different tasks. We utilize this metric to dynamically adjust the gate’s fusion rate of multi-level features for varied scenarios as
@@ 3.3 Training Strategy
We initially fine-tune the model to ensure multimodal feature alignment contributed to M33{3}3Adapter. For the synthesis of pertinent visuals, the extracted mapping feature $h_{align}$ and $h_{palign}$ serves as a conditional input during the denoising phase. Drawing parallels with Score Distillation Sampling (SDS) [35], we feed aligned text embeddings into SDXL’s frozen pre-trained UNet $\theta_{unet}$ to generate a score reflecting the target image distribution which can me mathematically formulated as:
@@
$\mathcal{L}_{ddpm}:=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1),\lambda}\left[||\epsilon-h_{unet}||^{2}\right]$ (5)
@@ 4.4 Qualitative Comparisons
As shown in Fig. 4, our pipeline generates high-resolution images in different contents. It is demonstrated that our efficient alignment methods adapt the prompts well, as described in the quantitative results. Our $M^{2}Chat$ without M3FT is compared with SDXL-base and SD1.5 for a fair comparison. Here, the generalization resolution is 1024$\times$1024. In conclusion, M2Chat is able to fit the prompt better than SD1.5. We provide more generation results in Appendix A.

### v3

@@ 3.1 Preliminary
We begin by detailing our instruction design process. We draw from an image-text dataset $\mathcal{D}:\{\mathcal{X,Y}\}$, containing pairs of images $\{x\}_{i=1}^{N}$ and their corresponding textual contexts $\{y\}_{i=1}^{N}$, where $N$ is the sample count. To construct the context $Y$, we adopt the principles of visual instruction tuning [21] and introduce an additional image token <|img|><|img|> to denote padding, alongside <|IC|><|IC|> to signal the start of an image caption. These tokens serve as markers to differentiate token types during the two-stage M3FT training phase. The context $y$ is organized as
@@ 3.2 Framework Architecture
where $l$ indicates the $l^{th}$ layer. However, aligning features from two modalities directly limits model creativity, particularly in multimodal conversation. Therefore, we introduce a multi-level feature fusion strategy that incorporates granular low-level visual primitive $f_{img}$ encoding layouts and textures into semantically rich deep high-level multimodal feature $h_{l}$ to obtain the final fused feature $f_{fus}$. A learnable gate modulates this fusion for adaptability. This method enriches image generation guidance, tailoring output to varied task demands. We differentiate answer tokens $t_{ans}$ and caption tokens $t_{cap}$ from VLM output $h_{l}$ with special tokens <|IC|><|IC|>, computing their cosine similarity, which acts as a task-specific flag signaling when answers and captions diverge or require consistency in different tasks. We utilize this metric to dynamically adjust the gate’s fusion rate of multi-level features for varied scenarios as
@@ 3.3 Training Strategy
We initially fine-tune the model to ensure multimodal feature alignment contributed to M3{3}Adapter. For the synthesis of pertinent visuals, the extracted mapping feature $h_{align}$ and $h_{palign}$ serves as a conditional input during the denoising phase. Drawing parallels with Score Distillation Sampling (SDS) [35], we feed aligned text embeddings into SDXL’s frozen pre-trained UNet $\theta_{unet}$ to generate a score reflecting the target image distribution which can me mathematically formulated as:
@@
[…]
@@ 4.4 Qualitative Comparisons
As shown in Fig. 4, our pipeline generates high-resolution images in different contents. It is demonstrated that our efficient alignment methods adapt the prompts well, as described in the quantitative results. Our $M^{2}Chat$ without M3FT is compared with SDXL-base and SD1.5 for a fair comparison. Here, the generalization resolution is $\times1024$. In conclusion, M2Chat is able to fit the prompt better than SD1.5. We provide more generation results in Appendix A.